# ConsDreamer: Advancing Multi-View Consistency for Zero-Shot Text-to-3D Generation

### v1

@@ III-B View Disentanglement Module
where $w_{2},w_{3}$ control the intensity of view injection. As illustrated in Fig.3, we implement the dynamic adjustment of injection intensity using Eq.(13) and Eq.(14), the model can more accurately “understand” the target view $\lambda_{n}$ specified by the rendering camera. Experimental results in Section IV. demonstrate that this approach effectively mitigates the multi-face Janus problem.
$\mathbf{V}_{{key}}^{prompt}\!\leftarrow\!\mathbf{V}_{{key}}^{prompt}\underbrace{\!-\!\sum^{\Omega}Proj_{\delta_{{key}}^{{view}}}\!\left(\!\mathbf{V}_{{key}}^{prompt}\!\right)\!}_{\text{\scriptsize Prior Elimination}}\underbrace{+\eta\cdot\delta_{{key}}^{{back}}}_{\text{\scriptsize ControlInjection}},\end{aligned}$
III-C Partial Order Loss for Cross-View Consistency
Figure 5: Partial order loss for cross-view consistency. The random camera views are distributed on the unit circle based on their azimuth angles. A reference camera (e.g., camera 0) is selected, and a red dotted line parallel to the horizontal axis is constructed from this point to mirror the reference to a symmetric position. The azimuthal distances between other camera views and the nearest reference are computed to determine the expected similarity relationships ($sim_{0,3}>sim_{0,2}>sim_{0,1}$). The right-hand side of the figure computes the actual similarity relationships using cosine similarity, and $\mathcal{L}_{P}$ evaluates the alignment between the actual and expected similarity relationships.
Figure 6: Impact of $\mathcal{L}_{P}$ on 3D content over iterations. Using “A squirrel eating a burger” as an example, $\mathcal{L}_{P}$ demonstrates a significant suppression effect on the multi-face Janus problem encountered during early training stages.
To achieve this, we explore the relationship between the multi-view rendered images. A random image is selected from all rendered images as the reference image, and the cosine similarity between the other images and the reference image is computed. As shown in Fig.5, the red camera represents the reference view selected from a series of randomly generated camera views $\Lambda$. It is evident that the closer the azimuth is to the reference azimuth, the higher the similarity score between the corresponding rendered image and the reference image. Furthermore, the scores gradually decrease symmetrically as the azimuth angle deviates from the reference. Therefore, we define a similarity-based partial order loss, where the specific ordering method is illustrated in Fig.5. A Cartesian coordinate system is established, and the random camera views are distributed on the unit circle based on their azimuth angles. One camera view is randomly chosen as the reference view, e.g., camera 0 in Fig.5, and a red dotted line parallel to the horizontal axis is constructed from this point. The reference camera is then mirrored to a symmetric point using this line. The azimuthal distances between other camera views and the nearest reference camera view are computed, determining the expected similarity partial order among the corresponding rendered images: $sim_{0,3}>sim_{0,2}>sim_{0,1}$. The similarity relationship between the 2D projections obtained from Eq.(8) must also adhere to this order. Accordingly, the cosine similarity between the corresponding rendered images is computed as the actual similarity and a similarity-based partial order loss $\mathcal{L}_{P}$ is employed to constrain the actual similarity to align with the expected similarity relationships. $\mathcal{L}_{P}$ is defined as:
$\mathcal{L}_{P}\!=\!\sum_{i=1}^{|\Lambda|-2}\!\max\!\left(0,{sim}\left(R^{i+1},R^{0}\right)\!-\!{sim}\left(R^{i},R^{0}\right)\right),$ (16)
@@ III-C Partial Order Loss for Cross-View Consistency
where $R^{i}$ is the 2D rendered image from the 3D content with the $i^{th}$ view in $\Lambda$, $R^{0}$ is the reference image and $\operatorname{sim}\left(R^{i},R^{0}\right)$ denotes the cosine similarity between $R^{i}$ and $R^{0}$. The $\mathcal{L}_{P}$ term can constrain the multi-face Janus problem early in training. As shown in Fig.6, using “A squirrel eating a burger” as an example, without the $\mathcal{L}_{P}$ constraint, the generated result exhibits an irreconcilable multi-face Janus problem. When $\mathcal{L}_{P}$ is applied, the generated 3D content shows significant improvement starting from the 1000th iteration.
Subsequently, we utilize a parameter update module based on DDIM and ISM, which originates from the methodologies outlined in LucidDreamer. Specifically, with a given prompt $c$ and the noisy latents $\boldsymbol{z}_{s}$ and $\boldsymbol{z}_{t}$ generated through DDIM inversion from $\boldsymbol{z}_{0}$ with the time interval $t-s$, the ISM loss is defined as:
$\mathcal{L}_{ISM}(\theta)\!=\!\mathbb{E}_{t,c}\!\left[\omega(t)\!\left\|\boldsymbol{\epsilon}_{\theta}\left(\boldsymbol{z}_{t},t,c\right)\!-\!\boldsymbol{\epsilon}_{\theta}\left(\boldsymbol{z}_{s},s,\emptyset\right)\right\|_{2}^{2}\right].$ (17)
For detailed information of ISM, please refer to [17]. The final loss function is then formalized as:
${Loss}=\mathcal{L}_{ISM}+\kappa\mathcal{L}_{P},$ (18)
where $\kappa$ is the weight coefficients of $\mathcal{L}_{P}$.
@@ IV Experiments
In this section, we provide a systematic evaluation of ConsDreamer through a dual approach integrating qualitative visual analysis and quantitative benchmarking. First, we conduct a comprehensive ablation study in Section IV-A to evaluate the effectiveness of each individual module. Next, Section IV-B presents comprehensive comparisons against seven SOTA methods to further highlight the advantages of our framework. Finally, we explore the improvement brought by VDM to T2I tasks in Section IV-C, further extending the evaluation. All experiments were conducted using a single NVIDIA RTX 4090 GPU and the Stable Diffusion v2.1
Figure 7: Qualitative results of ablation study. The first row shows the results without the intervention of VDM or $\mathcal{L}_{P}$. The second and third rows display the generated results after applying VDM or $\mathcal{L}_{P}$ respectively. The bottom row presents the performance with both modules jointly optimized. Each generated object is examined from different angles to visually assess its multi-view consistency.
base
model for distillation. Besides, we set the training process with 5,000 iterations and employed the pre-trained Point-E [31] to initialize the 3D Gaussian Splatting.
@@ IV-A Ablation Studies
TABLE I: Quantitative Results of Ablation Study on ConsDreamer
@@
[…]
Rank Score Rank Score Rank Score Rank Score Rank Rate Rank Rate
Janus issue 5 -0.770 1 0.340 5 0.086 5 0.050 5 100.0 5 100.0
w/o $\mathcal{L}_{p}$ w/o VDM 4 0.282 5 0.328 3 0.076 4 0.043 4 24.0 4 64.0
w/o VDM 3 0.294 4 0.332 3 0.076 3 0.042 3 15.3 3 52.7
w/o $\mathcal{L}_{p}$ 2 0.465 2 0.333 2 0.068 2 0.039 2 11.4 2 42.0
Ours 1 0.707 2 0.333 1 0.060 1 0.035 1 3.4 1 31.0
  • $\uparrow$
    Higher scores indicate better performance.
  • $\downarrow$
    Lower scores indicate better performance.
By eliminating view semantic bias in T2I mappings and incorporating explicit consistency constraints across multi-view images, our method significantly improves the consistency of generated content. As shown in Fig.7, both VDM and the perceptual loss $\mathcal{L}_{P}$ effectively reduce the occurrence of the multi-face Janus issue. However, certain inconsistency artifacts persist, as demonstrated in the second and third rows of Fig.7, including redundant Batman ears, additional Pikachu limbs, and extraneous nuclear reactors on Iron Man’s back. Notably, when these two modules are combined, the proposed method achieves highly consistent 3D content, surpassing the performance of using either module individually. The results in the bottom row of Fig.7 demonstrate that our approach not only mitigates the multi-face Janus issue but also effectively prevents component redundancy.
Figure 10: Qualitative comparisons of 3D content generated by different methods. This figure presents a comparison of our method with seven SOTA text-to-3D generation approaches across various objects, evaluated from five viewpoints (0°, 45°, 90°, 135°, and 180°) for quality and consistency. When generating Super Saiyan characters, DreamFusion, Sherpa3D, LucidDreamer, ScaleDreamer, DreamScene, and GCS-BEG produced redundant limbs. For Iron Man, Sherpa3D, LucidDreamer, ScaleDreamer, and DreamScene incorrectly generated a nuclear reactor on the back. In the case of Batman, Sherpa3D, LucidDreamer, DreamScene, and GCS-BEG generated three ears. The results demonstrate that our proposed ConsDreamer outperforms SOTA methods in terms of both generation quality and multi-view consistency.
TABLE II: Quantitative Comparison with SOTA Methods
[…]
Rank Score Rank Score Rank Score Rank Score Rank Score Rank Rate Rank Rate
DreamFusion [14] 8 3.48 8 4.37 8 3.93 7 -0.289 6 0.307 8 60.0 8 86.7
Debiased-SDS [15] 7 4.19 7 4.93 7 4.56 8 -0.563 8 0.282 4 46.7 2 60.0
Sherpa3D [22] 6 5.42 6 5.82 6 5.62 6 -0.161 7 0.291 7 56.3 7 75.0
LucidDreamer [17] 4 6.76 5 6.31 4 6.54 4 0.285 4 0.328 2 24.0 3 64.0
ScaleDreamer [24] 2 7.82 4 6.33 3 7.08 2 0.371 1 0.333 5 47.3 6 69.2
DreamScene [23] 5 6.25 3 6.90 5 6.58 5 0.201 5 0.321 6 48.3 5 66.7
GCS-BEG [21] 3 7.03 2 7.65 2 7.34 3 0.333 3 0.316 3 33.3 4 64.3
Ours 1 7.85 1 8.05 1 7.95 1 0.707 1 0.333 1 3.4 1 31.0
  • $\uparrow$
    Higher scores indicate better performance.
  • $\downarrow$
    Lower scores indicate better performance.
2) Quantitative Results: Qualitative results visually demonstrate the effectiveness of the proposed framework. We employed conventional evaluation metrics in text-to-3D generation, including ImageReward [50], OpenCLIP-L/14 [51], and A-LPIPS [52], to objectively assess the effectiveness of our framework. ImageReward measures the consistency between generated results and text descriptions, OpenCLIP-L/14 provides enhanced text-image alignment capabilities to capture fine-grained semantic information, and A-LPIPS evaluates the visual quality of generated images, focusing on perceptual similarity. Table I shows that each module of our method outperforms the scenario without intervention in various aspects, and the joint optimization of both modules achieves even better performance. Additionally, we evaluated the percentage of generated results exhibiting Inconsistency, denoted as $f_{inc}$. We broadened the traditional definition of the multi-face Janus problem to encompass a wider range of inconsistencies, such as extra limbs and other implausible artifacts. Fig. 8 provides illustrative examples of Inconsistency. For a comprehensive evaluation, we selected a diverse set of prompts involving objects with distinct front-back differences (e.g., portraits, animals, and vehicles). The quantitative results are summarized in Table I, where the frequency of the multi-face Janus problem is denoted as $f_{mf}$. The results indicate that any of our proposed components significantly improves the consistency of the generated content. Specifically, applying $\mathcal{L}_{p}$ achieves significant improvements, reducing the multi-face problem and enhancing overall consistency by 36.25% and 18.75%, respectively, compared to the scenario without intervention. Similarly, applying VDM demonstrates further enhancements, with improvements of 52.5% and 34.38%, respectively. When both modules are jointly applied, the performance gains are substantially amplified, reaching 85.83% and 51.56%, respectively.
@@
However, we chose those generation results with the multi-face Janus problem and observed that the OpenCLIP-L/14 metric yields higher scores in Table I. Fig.9 presents the OpenCLIP similarity distribution between prompts and rendered results for different methods, including samples exhibiting the multi-face Janus problem for comparison. The similarity scores generally decrease progressively from the front view toward both ends, with the lowest scores observed for the back view, which aligns with intuition, as rendered results from the back view typically contain less information, leading to lower similarity between the prompt and the image for that view. However, when the generated object exhibits the multi-face Janus problem, features of prior preference views appear across multiple angles, resulting in abnormally high similarity scores for those views. This observation prompts us to explore more suitable ablation study methods. To more accurately evaluate the consistency of 3D-generated content, analysing the distribution of similarity scores across views is more effective than relying solely on the average similarity score. The results demonstrate that our method produces a similarity score distribution that aligns more closely with the ideal one.
IV-B Comparison with SOTA Methods
1) Qualitative Results: Fig.2 showcases the vivid and diverse text-to-3D results generated by our method. Subsequently, a qualitative comparison is conducted against seven SOTA text-to-3D generation methods. To better assess the quality and view consistency of the generated models, five evenly distributed viewpoints over a 180° azimuth rotation are evaluated. As demonstrated in Fig.10, our method substantially alleviates the multi-face Janus issue prevalent in existing approaches.
2) Quantitative Results: We selected ImageReward, OpenCLIP-L/14, and the frequency of Inconsistency as objective evaluation metrics. As shown in Table II, our method achieves the highest scores across all metrics, particularly in the rate of the multi-face Janus issue, where our method outperforms the average by approximately 90%.
We conducted a user study involving 105 participants to compare with other SOTA methods as subjective evaluation results. Using a 1 to 10 scale, participants assessed the models based on two key criteria: “Generation Quality” and “ 3D Consistency.” To ensure a fair and unbiased comparison, all samples used in the study were randomly selected, with no cherry-picking to favor our approach. As shown in Table II, our method achieves the highest average scores across both evaluation metrics compared to the seven competing methods. These results demonstrate that the 3D content generated by our approach is consistently more appealing and exhibits superior view consistency.

### v2

@@ III-C Partial Order Loss for Cross-View Consistency
Rank Score Rank Score Rank Score Rank Score Rank Score Rank Score
Janus Issue 7 -0.770 1 0.340 7 0.086 7 0.050 7 100.0 7 100.0
Baseline 6 -0.637 2 0.338 6 0.079 6 0.047 6 85.7 6 97.1
Baseline + $\mathcal{L}_{\text{P}}$ 5 0.274 7 0.327 3 0.074 3 0.042 5 37.1 4 60.0
Baseline + Perp-Neg 4 0.285 6 0.328 4 0.076 5 0.043 4 24.0 5 64.0
Baseline + Perp-Neg + $\mathcal{L}_{\text{P}}$ 3 0.294 5 0.332 4 0.076 3 0.042 3 15.3 3 52.7
Baseline + VDM 2 0.465 3 0.333 2 0.068 2 0.039 2 11.4 2 42.0
Baseline + VDM + $\mathcal{L}_{\text{P}}$ (Ours) 1 0.707 3 0.333 1 0.060 1 0.035 1 3.4 1 31.0
@@
where $F^{i}$ is the encoded feature of the rendered image corresponding to the $i^{th}$ view in $\Lambda$, $F^{0}$ is the reference feature and $\operatorname{sim}\left(F^{i},F^{0}\right)$ denotes the cosine similarity between $F^{i}$ and $F^{0}$. The $\mathcal{L}_{\text{P}}$ term can constrain the multi-face Janus problem early in training. As shown in Fig.7, using “A squirrel eating a burger” as an example, without the $\mathcal{L}_{\text{P}}$ constraint, the generated result exhibits an irreconcilable multi-face Janus problem. When $\mathcal{L}_{\text{P}}$ is applied, the generated 3D content shows significant improvement starting from the 1000th iteration.
Our plug-and-play method integrates into multiple baseline frameworks with diverse score distillation strategies, unifying their score distillation losses into a consistent formulation $\mathcal{L}_{\text{score}}$. The final loss function is then formalised as:
${Loss}=\mathcal{L}_{\text{score}}+\kappa\mathcal{L}_{\text{P}},$ (18)
where $\kappa$ is the weight coefficient of $\mathcal{L}_{\text{P}}$.
@@ IV Experiments
In this section, we provide a systematic evaluation of ConsDreamer through a dual approach integrating qualitative visual analysis and quantitative benchmarking. First, we conduct a comprehensive ablation study in Section IV-A to evaluate the effectiveness of each individual module. Furthermore, Perp-Neg is additionally considered as an ablation variable to ensure a fair and comprehensive comparison. Next, in Section IV-B, ConsDreamer is integrated into six SOTA baselines encompassing diverse 3D representations (NeRF or 3DGS) and various score distillation strategies to validate the versatility of our method. Finally, we explore the improvement brought by VDM to T2I tasks in Section IV-D, further extending the evaluation. All experiments were conducted using the Stable Diffusion v2.1 base model for distillation.
Figure 9: Ablation on weight coefficient $\kappa$ of $\mathcal{L}_{\text{P}}$. $\kappa$ = 0.6 yields the highest visual quality: excessively large values lead to over-smoothing and misalignment with textual prompts (e.g., “white hair”), while overly small values cause view inconsistency in the generated results, resulting in poor fitting of the target object.
@@ IV-A Ablation Studies
By eliminating view semantic bias in T2I mappings and incorporating explicit consistency constraints across multi-view images, our method significantly improves the consistency of generated content. As illustrated in Fig.8, the baseline model without any consistency strategy suffers from severe multi-face Janus artefacts, which are characterised by additional limbs and redundant faces of Pikachu. When the perceptual loss $\mathcal{L}_{\text{P}}$ is applied, the extra limbs and ears are effectively suppressed. To investigate the performance gain of VDM over Perp-Neg, we also treat Perp-Neg as an ablation variable, and the visualisation results demonstrate a notable improvement with VDM achieving more consistent generation results than Perp-Neg. However, certain inconsistent artefacts, such as redundant limbs still persist. The model yields the optimal performance in terms of both consistency and visual quality when VDM is combined with $\mathcal{L}_{\text{P}}$. Furthermore, as shown in Fig. 9, we conducted an ablation study on the weight coefficient $\kappa$ of the partial order loss $\mathcal{L}_{\text{P}}$.
2) Quantitative Results: Qualitative results visually demonstrate the effectiveness of the proposed method. We employed conventional evaluation metrics in text-to-3D generation, including ImageReward [NEURIPS2023_33646ef0-imagereward], OpenCLIP-L/14 [ramesh2022hierarchical-openclip], and A-LPIPS [Zhang_2018_CVPR-alpips], to objectively assess the effectiveness of our method. ImageReward measures the consistency between generated results and text descriptions, OpenCLIP-L/14 provides enhanced text-image alignment capabilities to capture fine-grained semantic information, and A-LPIPS evaluates the visual quality of generated images, focusing on perceptual similarity. Table II shows that each module of our method outperforms the scenario without intervention in various aspects, and the joint optimisation of both modules achieves even better performance. Additionally, we evaluated the percentage of generated results exhibiting Inconsistency, denoted as $\text{f}_{\text{inc}}$. We broadened the traditional definition of the multi-face Janus problem to encompass a wider range of inconsistencies, such as extra limbs and other implausible artefacts. Fig. 10 provides illustrative examples of Inconsistency. For a comprehensive evaluation, we selected a diverse set of prompts involving objects with distinct front-back differences (e.g., portraits, animals, and vehicles). The quantitative results are summarised in Table II, where the frequency of the multi-face Janus problem is denoted as $\text{f}_{\text{mf}}$. The results indicate that each proposed component significantly improves the consistency of the generated content. Notably, when directly comparing VDM against Perp-Neg, VDM demonstrates substantially superior performance, achieving an additional 52.5% reduction in $\text{f}_{\text{mf}}$ (from 24.0% to 11.4%) and a 34.4% reduction in $\text{f}_{\text{inc}}$ (from 64.0% to 42.0%). This significant improvement validates VDM’s effectiveness in addressing prior view biases. When combined with $\mathcal{L}_{\text{P}}$, the performance gains are further amplified. Our complete method (VDM + $\mathcal{L}_{\text{P}}$) achieves the best overall performance with $\text{f}_{\text{mf}}$ of only 3.4% and $\text{f}_{\text{inc}}$ of 31.0%, representing reductions of 96.0% and 68.1% compared to the baseline, respectively. Compared to Perp-Neg + $\mathcal{L}_{\text{P}}$, our method further reduces $\text{f}_{\text{mf}}$ by 77.8% and $\text{f}_{\text{inc}}$ by 41.2%, demonstrating the synergistic effects of VDM and $\mathcal{L}_{\text{P}}$ in effectively mitigating the multi-face Janus problem.
Figure 12: Qualitative text-to-3D generation results of our method. We integrate ConsDreamer into six SOTA baselines spanning NeRF/3DGS representations and diverse score distillation strategies, and these results demonstrate that our proposed method is visually superior to the baselines.
However, we chose those generation results with the multi-face Janus problem and observed that the OpenCLIP-L/14 metric yields higher scores in Table II. Fig.11 presents the OpenCLIP similarity distribution between prompts and rendered results for different methods, including samples exhibiting the multi-face Janus problem for comparison. The similarity scores generally decrease progressively from the front view toward both ends, with the lowest scores observed for the back view, which aligns with intuition, as rendered results from the back view typically contain less information, leading to lower similarity between the prompt and the image for that view. However, when the generated object exhibits the multi-face Janus problem, features of prior preference views appear across multiple angles, resulting in abnormally high similarity scores for those views. This observation prompts us to explore more suitable ablation study methods. To more accurately evaluate the consistency of 3D-generated content, analysing the distribution of similarity scores across views is more effective than relying solely on the average similarity score. The results demonstrate that our method produces a similarity score distribution that aligns more closely with the ideal one.
TABLE III: Quantitative Results of Integrating ConsDreamer into Various Baselines
Methods 3D Representation Score Distillation Consistency Strategy User Study ImageReward $\uparrow$ OpenCLIP $\uparrow$ Frequency of Inconsistency Time (min) VRAM (GiB)
@@
DreamFusion NeRF SDS \ 3.48 4.37 3.93 $0.289\pm 0.81$ $0.307\pm 0.025$ 60.0 86.7 $\sim$ 40 6.2
+ Ours NeRF SDS ConsDreamer 5.45 5.58 5.52 $\textbf{0.103}\pm 0.79$ $\textbf{0.317}\pm 0.023$ 26.7 60.0 $\sim$ 42 6.3
SDI NeRF SDI Perp-Neg 6.33 6.89 6.61 $0.297\pm 0.61$ $0.328\pm 0.024$ 33.3 61.9 $\sim$ 119 39.2
+ Ours NeRF SDI ConsDreamer 6.51 7.03 6.77 $\textbf{0.175}\pm 0.59$ $\mathbf{0.331}\pm 0.021$ 27.0 40.5 $\sim$ 116 39.0
DreamScene 3DGS FPS Perp-Neg 6.25 6.90 6.58 $0.201\pm 0.75$ $0.321\pm 0.027$ 48.3 66.7 $\sim$ 30 13.6
+ Ours 3DGS FPS ConsDreamer 6.93 7.45 7.19 $\mathbf{0.531}\pm 0.52$ $\textbf{0.324}\pm 0.027$ 28.3 45.0 $\sim$ 30 13.6
LucidDreamer 3DGS ISM Perp-Neg 6.76 6.31 6.54 $0.285\pm 0.69$ $0.328\pm 0.019$ 24.0 64.0 $\sim$ 55 22.1
+ Ours 3DGS ISM ConsDreamer 7.85 8.05 7.95 $\mathbf{0.707}\pm 0.41$ $\mathbf{0.333}\pm 0.015$ 3.4 31.0 $\sim$ 55 22.0
GCS-BEG 3DGS GCS Perp-Neg 7.03 7.65 7.34 $0.333\pm 0.83$ $0.316\pm 0.031$ 33.3 64.3 $\sim$ 41 16.4
+ Ours 3DGS GCS ConsDreamer 7.97 8.16 8.07 $\mathbf{0.814}\pm 0.51$ $\textbf{0.327}\pm 0.035$ 7.1 35.7 $\sim$ 41 16.3
GE3D 3DGS GE3D Perp-Neg 6.58 6.93 6.76 $0.297\pm 0.47$ $0.324\pm 0.020$ 34.3 65.7 $\sim$ 109 12.5
+ Ours 3DGS GE3D ConsDreamer 7.01 7.64 7.33 $\mathbf{0.447}\pm 0.44$ $\textbf{0.324}\pm 0.019$ 17.1 36.7 $\sim$ 109 12.5
IV-B Integration with SOTA Methods
1) Qualitative Results: Fig.2 presents vivid and diverse text-to-3D generation results produced by our method. Furthermore, we integrate ConsDreamer into a range of advanced 3D generation frameworks, which cover different 3D representation paradigms and various score distillation strategies. As demonstrated in Fig.12, our method substantially alleviates the multi-face Janus issue prevalent in existing approaches.
@@ IV-B Integration with SOTA Methods
2) Quantitative Results: We selected ImageReward, OpenCLIP-L/14, and inconsistency frequency as objective evaluation metrics. As shown in Table III, ConsDreamer delivers more substantial gains over the baseline than PerpNeg, attaining state-of-the-art performance across all metrics. Specifically, in the multi-face Janus issue rate, our method outperforms PerpNeg-based baselines by approximately 52.3%. Additionally, we report the mean runtime and GPU VRAM usage. For GPU VRAM usage, we compute the average of the maximum GPU memory consumption across multiple runs. Table III demonstrates that our method is lightweight, as it does not compromise GPU memory or runtime for consistent generation.
We conducted a user study involving 145 participants to compare with other SOTA methods as subjective evaluation results. Using a 1 to 10 scale, participants assessed the models based on two key criteria: “Generation Quality” and “ 3D Consistency.” To ensure a fair and unbiased comparison, all samples used in the study were randomly selected, with no cherry-picking to favour our approach. As shown in Table III, our method achieves the highest average scores across both evaluation metrics compared to baselines. These results demonstrate that the 3D content generated by our approach is consistently more appealing and exhibits superior view consistency.